# Evidence-Focused Fact Summarization for Knowledge-Augmented Zero-Shot Question Answering

### v1

@@ 4.3 Effectiveness of Clear Evidence (RQ2)
Comparison with other baselines. To examine the effectiveness of EFSum in terms of evidence clarity, we evaluate the LLM’s zero-shot QA accuracy when the number of facts is and . In Table 3, EFSum shows the highest performance across most datasets, showing consistently leading results for both values. Nevertheless, we notice the presence of uncontrolled model inclination. For example, in Tables 1 and 3, KAPING consistently exhibits the highest performance when Llama2-7B-Chat is utilized as the QA model for the WebQSP dataset. This might be due to the model’s inclination towards a specific knowledge format; irrespective of a verbalization method’s overall superiority, certain QA models may better understand a specific knowledge format in particular datasets. Excluding these uncontrollable model inclinations, EFSum generally shows dominant results than other baselines.
Robustness of EFSum across various . To assess the robustness of EFSum, we increase the retrieved number of facts from 10 to 150 and measure the performance of each baseline at the answer-level and summary-level. At the answer-level, we evaluate the extent to which the answers are contained within the responses generated by the QA model using summaries. In case of the summary-level, we evaluate whether the answer is included within the verbalized knowledge produced by the method. In Table 4, across both datasets and the majority of K values, EFSum consistently outperforms others. Notably, in terms of summary-level accuracy, EFSum also demonstrates superior performance, surpassing all methods except for the simple linearization approach, KAPING, which inherently achieves a summary level accuracy of 1. Although our method do not achieve a summary-level accuracy of 1, its superior performance compared to other baselines in terms of answer-level accuracy suggests that our method is highly effective in aiding question-answering tasks.
@@ 4.4 Effect of Preference Alignment (RQ3)
Question: Where was george washington carver from? Answer: Diamond |
KAPING: (George Washington Carver, place of birth, Diamond),(George Washington Carver, given name, George),(George Washington Carver, occupation, biologist),(George Washington Carver, interested in, botany),(George Washington Carver, occupation, university teacher),(George Washington Carver, field of work, pedagogy),(George Washington Carver, relative, Moses Carver),(George Washington Carver, place of birth, United States of America),(George Washington Carver, residence, Tuskegee),(George Washington Carver, occupation, inventor) |
Rewrite: George Washington Carver is associated with the Commons category "George Washington Carver". George Washington Carver was a botanist. George Washington Carver was an inventor. George Washington Carver was a university teacher. George Washington Carver was a mycologist. George Washington Carver was a chemist. George Washington Carver was a biologist. George Washington Carver was a citizen of the United States of America. George Washington Carver died on January 5, 1943. George Washington Carver is a human. |
EFSum: George Washington Carver, the renowned biologist, was born in Diamond, United States of America. He took great interest in botany and became a university teacher, focusing on pedagogy. Additionally, Carver contributed significantly as an inventor. His relative, Moses Carver, also had a close association with him. Carver later resided in Tuskegee. |
Helpfulness and faithfulness. We examine the helpfulness and faithfulness of the verbalized facts generated by each verbalization method.555We opt not to include KAPING in the plot, as its linear verbalization does not alter the content of the facts. Note that helpfulness is calculated through summary-level accuracy, while faithfulness is calculated as 1 - (hallucination occurrence rate). That is, a method with a lower rate of hallucination occurrences possesses higher faithfulness. Figure 5 illustrates that EFSum achieves superior helpfulness and faithfulness compared to other baselines, including EFSum. In other words, EFSum is less likely to generate summaries of hallucination and is capable of incorporating more correct answers into the summaries compared to other baselines. Moreover, it validates that using the two filters in the broad-to-specific paraphrasing process definitively aids in improving both faithfulness and helpfulness.
Case study. Finally, we present an example of verbalized facts generated by each method. In Table 4, EFSum emphasizes evidence necessary to answer the question (i.e., was born in Diamond) at the beginning, while excluding irrelevant details for brevity, as opposed to other baselines. Duplicated relations (i.e., occupation) are removed, and considering the relationships between other facts, it is converted into more natural sentences.
@@ 5.2 Knowledge Graph Question Answering
Studies exploring the integration of additional knowledge extracted from KGs, represented as subgraphs, are divided into two main approaches. The first approach utilizes semantic parsing techniques Bao et al. (2016); Luo et al. (2018), which enable the extraction of executable queries from the KG by using contextual information as a parsing reference. Alternatively, the information retrieval approach involves encoding assimilated information using techniques like Graph Neural Networks (GNNs). Several recent studies Yasunaga et al. (2022a); Zhang et al. (2022); Yasunaga et al. (2022) propose a learning framework that combines GNNs and LMs, allowing concurrent utilization of textual data and KG. In contrast to the conventional KGQA approaches that aim to directly identify the answer entity within KGs, in this work, we mainly focus on how the KG-retrieved facts can be effectively utilized within the zero-shot QA capability of LLMs.
@@ 7 Limitation
Secondly, there are cases where the tendency of LLM to favor a certain fact verbalization method becomes overwhelming and difficult to manage. As demonstrated in Tables 2 and 3, with the example of Llama2-7B-Chat on the WebQSP dataset, it has been observed that certain models may have a inclination for specific knowledge formats in particular datasets. Consequently, while our summarizer generally demonstrates good performance, controlling performance may become challenging when a specific model has a strong inclination towards a particular knowledge format.
Lastly, it is important to note that the performance of our proposed summarizer can be influenced by the performance of the retriever. As can be seen in Table 2, using a better retriever can lead to higher performance. While the off-the-shelf model we used (i.e., MPNet) demonstrates excellent retrieving capabilities based on the semantic similarity between questions and facts, it’s difficult to assert that it is a flawless retriever. Therefore, we are currently conducting further research to propose not only a more powerful summarizer but also a more flawless retrieving method simultaneously.
@@ A.1 LLMs for Zero-Shot QA
Flan-T5 is a variant model grounded in the T5 architecture. T5, an encoder-decoder model, is trained on a text-to-text dataset featuring a diverse array of objectives. Flan-T5 represents an evolution of T5 through the process of instruction tuning, enhancing its performance by aligning it more closely with specific instructional contexts. We employ Flan-T5-XL for our QA model
@@
GPT-3.5-turbo, developed by OpenAI, stands as a prominent closed-source model renowned for its high performance, making it particularly well-suited for application in QA models. Due to GPT-3.5-turbo being closed-source, we utilize the API provided by OpenAI666https://openai.com/api/.
@@ B.1 Datasets
WebQuestionsSP (WebQSP) Berant et al. (2013); Yih et al. (2016) Semantic Parses Dataset, abbreviated to WebQSP, is a KGQA dataset providing SPARQL queries which allows a direct retrieval from Freebase KG. Due to the cessation of updates to Freebase we adopt Wikidata as the foundational KG. For convenience, we skip entity linking and utilize the WebQSP-WD Sorokin and Gurevych (2018) dataset, wherein all questions are linked to entities within Wikidata.
Mintaka Sen et al. (2022) A KGQA dataset collected from Wikidata, with 8 complexity types. As mentioned in 4.1, we find that most types of complexity are not suitable for the generative KGQA task. For a more accurate evaluation, similar to previous work Razzhigaev et al. (2023), we decide to use only the generic type questions from Mintaka. Out of 4000 questions in the Mintaka test set, we only use 795 generic type questions for evaluation.
@@ C.1 Case Study
Question What did st augustine do? 777This example is from WebQSP dataset
Answers [“Philosopher”, “Physician”, “Writer”]
KAPING
(Augustine of Hippo, occupation, Catholic bishop), (Augustine of Hippo, occupation, titular bishop), (Augustine of Hippo, occupation, theologian), (Augustine of Hippo, occupation, historian), (Augustine of Hippo, occupation, Catholic priest), (Augustine of Hippo, position held, diocesan bishop), (Augustine of Hippo, occupation, autobiographer), (Augustine of Hippo, occupation, poet), (Augustine of Hippo, date of birth, 0354-11-13), (Augustine of Hippo, occupation, preacher), (Augustine of Hippo, place of birth, Thagaste), (Augustine of Hippo, feast day, August 28), (Augustine of Hippo, date of birth, 0354-01-01), (Augustine of Hippo, consecrator, Megalius of Calama), (Augustine of Hippo, date of death, 0430-08-28), (Augustine of Hippo, occupation, presbyter), (Augustine of Hippo, feast day, June 15), (Augustine of Hippo, described by source, Church Encyclopedia), (Augustine of Hippo, father, Patricius), (Augustine of Hippo, occupation, writer), (Augustine of Hippo, described by source, Catholic Encyclopedia), (Saint Augustine, Bishop of Hippo (354-430) in Meditation, location, Kingston Lacy), (Augustine of Hippo, date of death, 0430-01-01), (Augustine of Hippo, child, Adeodatus), (Augustine of Hippo, notable work, The City of God), (Augustine of Hippo, lifestyle, mysticism), (Augustine of Hippo, occupation, philosopher), (Augustine of Hippo, honorific prefix, Doctor of the Church), (Augustine of Hippo, ethnic group, Berber), (Augustine of Hippo, place of death, Hippo Regius)
KG2Text
augustine of hippo was born on the 13th of november , 0354 . 01 . 18 . he was a catholic priest and a theologian . he was a poet , a musician and a musician . he was a member of the roman catholic church and was a member of the roman catholic church . he was a member of the roman catholic church and was a member of the roman catholic order . he was a member of the roman catholic order and was a member of the roman catholic order . he was
Rewrite
Augustine of Hippo is categorized under the main category “Category:Augustine of Hippo”. Augustine of Hippo is associated with the Commons category called “Augustinus”. Augustine of Hippo was born on November 13, 0354. Augustine of Hippo was born on January 1, 354. Augustine of Hippo died on August 28, 430. Augustine of Hippo died in 430. Augustine of Hippo is an instance of a human. Augustine of Hippo died in Hippo Regius. Augustine of Hippo was born in Thagaste. Augustine of Hippo’s feast day is celebrated on August 28. Augustine of Hippo’s feast day is on June 15. Augustine of Hippo is considered a saint. Augustine of Hippo was a philosopher. Augustine of Hippo was a theologian. Augustine of Hippo was an autobiographer. Augustine of Hippo was a music theorist. Augustine of Hippo was a preacher. Augustine of Hippo was a historian. Augustine of Hippo was a poet. Augustine of Hippo was a presbyter. Augustine of Hippo was a writer. Augustine of Hippo was a Catholic priest. Augustine of Hippo held the occupation of titular bishop. Augustine of Hippo was a Catholic bishop. Augustine of Hippo held the position of diocesan bishop. Augustine of Hippo was a citizen of Ancient Rome. Augustine of Hippo spoke, wrote, or signed in Latin. Augustine of Hippo spoke, wrote, or signed in the Punic language. Augustine of Hippo spoke, wrote, or signed in the Berber language. Augustine of Hippo, also known as Augustinus, is the creator of the Commons Creator page.
EFSum
Augustine of Hippo was a Catholic bishop and a theologian. He was also a historian, preacher, and poet. He was born in Thagaste in 354 and died in Hippo Regius in 430. He was consecrated as a bishop and served as the diocesan bishop of Hippo Regius. Augustine is well known for his notable work “The City of God” and is considered a Doctor of the Church. He was a Berber ethnicity and had a mystical lifestyle.

### v2

@@ 4.3 Effectiveness of Clear Evidence (RQ2)
Comparison with other baselines. To examine the effectiveness of EFSum in terms of evidence clarity, we evaluate the LLM’s zero-shot QA accuracy when the number of facts is and . In Table 3, EFSum shows the highest performance across most datasets, showing consistently leading results for both values. Nevertheless, we notice the presence of uncontrolled model inclination. For example, KAPING consistently exhibits the highest performance when Llama2-7B-Chat is utilized as the QA model for the WebQSP dataset. This might be due to the model’s inclination towards a specific knowledge format.
Robustness of EFSum across various . To assess the robustness of EFSum, we increase the retrieved number of facts from 10 to 150 and measure the performance of each baseline at the answer-level and summary-level. At the answer-level, we evaluate the extent to which the answers are contained within the responses generated by the QA model using summaries. In case of the summary-level, we evaluate whether the answer is included within the verbalized knowledge produced by the method. In Figure 4, across both datasets and the majority of K values, EFSum consistently outperforms, especially in answer-level accuracy, except for KAPING, which inherently achieves a summary level accuracy of 1 by simple linearization of facts.
@@ 4.4 Effect of Preference Alignment (RQ3)
Helpfulness and faithfulness. We examine the helpfulness and faithfulness of the verbalized facts generated by each verbalization method.444We opt not to include KAPING in the plot, as its linear verbalization does not alter the content of the facts. Note that helpfulness is calculated through summary-level accuracy, while faithfulness is calculated as 1 - (hallucination occurrence rate). That is, a method with a lower rate of hallucination occurrences possesses higher faithfulness. Figure 5 illustrates that EFSum achieves superior helpfulness and faithfulness compared to other baselines, including EFSum. In other words, EFSum is less likely to generate summaries of hallucination and is capable of incorporating more correct answers into the summaries compared to other baselines. Moreover, it validates that using the two filters in the broad-to-specific paraphrasing process definitively aids in improving both faithfulness and helpfulness.
@@ 5.2 Knowledge Graph Question Answering
Studies exploring the integration of additional knowledge extracted from KGs, represented as subgraphs, are divided into two main approaches. The first approach utilizes semantic parsing techniques Bao et al. (2016); Luo et al. (2018), which enable the extraction of executable queries from the KG by using contextual information as a parsing reference. Alternatively, the information retrieval approach involves encoding assimilated information using techniques like Graph Neural Networks (GNNs). Several recent studies Yasunaga et al. (2022a, b); Zhang et al. (2022) propose a learning framework that combines GNNs and LMs, allowing concurrent utilization of textual data and KG. In contrast to the conventional KGQA approaches that aim to directly identify the answer entity within KGs, in this work, we mainly focus on how the KG-retrieved facts can be effectively utilized within the zero-shot QA capability of LLMs.
@@ 7 Limitation
Secondly, there are cases where the tendency of LLMs to favor a certain fact verbalization method becomes overwhelming and difficult to manage. As demonstrated in Tables 2 and 3, with the example of Llama2-7B-Chat on the WebQSP dataset, it has been observed that certain models may have a inclination for specific knowledge formats in particular datasets. Consequently, while our summarizer generally demonstrates good performance, controlling performance may become challenging when a specific model has a strong inclination towards a particular knowledge format.
Lastly, it is important to note that the performance of our proposed summarizer can be influenced by the performance of the retriever. As can be seen in Table 2, using a better retriever can lead to higher performance. While the off-the-shelf model we used (i.e. MPNet) demonstrates retrieving capabilities based on the semantic similarity between questions and facts, it’s difficult to assert that it is a flawless retriever. For example, in a 2-hop dataset(i.e. Mintaka), it tends to retrieve 1-hop neighbors more than 2-hops even if it is irrelevant. This is because an answer entity in 2-hop neighbors is unseen in given question, so that a retrieval model may measure question entities in 1-hop neighbors more similar. Therefore, we are currently conducting further research to propose not only a more powerful summarizer but also a more flawless retrieving method simultaneously.
@@ A.1 LLMs for Zero-Shot QA
Flan-T5 is a variant model grounded in the T5 architecture. T5, an encoder-decoder model, is trained on a text-to-text dataset featuring a diverse array of objectives. Flan-T5 represents an evolution of T5 through the process of instruction tuning, enhancing its performance by aligning it more closely with specific instructional contexts. We employ Flan-T5-XL for our QA model.
@@
GPT-3.5-turbo, developed by OpenAI, stands as a prominent closed-source model renowned for its high performance, making it particularly well-suited for application in QA models. Due to GPT-3.5-turbo being closed-source, we utilize the API provided by OpenAI555https://openai.com/api/.
@@ B.1 Datasets
Mintaka Sen et al. (2022) A KGQA dataset collected from Wikidata, with 8 complexity types, including ‘Count’, ‘Comparative’, ‘Superlative’, ‘Ordinal’, ‘Multi-hop’, ‘Intersection’, ‘Difference’, ‘Yes/No’, ‘Generic’. Question-Answer pairs are collected from Wikidata entities. This dataset is multilingual, and we use English datasets in this work.
@@ C.1 Case Study
Question: Who is Emilio Estevez’s father? Answer: Martin Sheen |
KAPING: (Emilio Estevez’s, family name, Estévez), (Emilio Estevez’s, given name, Emilio), (Emilio Estevez’s, father, Martin Sheen), (Emilio Estevez’s, occupation, actor), (Joe Estevez, relative, Emilio Estevez), (Ramón Estévez, sibling, Emilio Estevez), (Emilio Estevez’s, relative, Joe Estevez), (Ramón Estévez, father, Martin Sheen), (Martin Sheen, child, Emilio Estevez), (Emilio Estevez’s, date of birth, +1962-05-12T00:00:00Z), (Joe Estevez, relative, Ramón Estévez), (Emilio Estevez’s, birth name, Emilio Estevez Sheen), (Emilio Estevez’s, sibling, Ramón Estévez), (Ramón Estévez, relative, Joe Estevez), (Ramón Estévez, family name, Estévez), (Emilio Estevez’s, occupation, film producer), (Renée Estévez, sibling, Emilio Estevez), (Charlie Sheen, sibling, Emilio Estevez), (Janet Sheen, child, Emilio Estevez), (Martin Sheen, child, Ramón Estévez), (Joe Estevez, family name, Estevez), (Martin Sheen, family name, Estévez), (Emilio Estevez’s, described by source, Obalky knih.cz), (Ramón Estévez, occupation, television actor), (Emilio Estevez’s, occupation, film director), (Emilio Estevez’s, country of citizenship, United States of America), (Emilio Estevez’s, occupation, writer), (Estévez, said to be the same as, Estevez), (Emilio Estevez’s, sibling, Renée Estévez), (Emilio Estevez’s, place of birth, New York City) |
KG2Text the family name of emilio estevez is estévez . he was born in new york city and is from the united states . he is the father of two children , both of whom are children . he is the actor , writer and film director . he is the brother of charlie sheen and his sister is renée . |
Rewrite: Emilio Estevez’s family name is Estévez. Emilio Estevez’s given name is Emilio. Emilio Estevez’s father is Martin Sheen. Emilio Estevez is an actor. Joe Estevez is a relative of Emilio Estevez. Ramón Estévez is a sibling of Emilio Estevez. Emilio Estevez is related to Joe Estevez as they are relatives. Ramón Estévez is the father of Martin Sheen. Martin Sheen is the father of Emilio Estevez. Joe Estevez is a relative of Ramón Estévez. Emilio Estevez has a sibling named Ramón Estévez. Ramón Estévez is a relative of Joe Estevez. Ramón Estévez’s family name is Estévez. Emilio Estevez is a film producer. Renée Estévez is the sibling of Emilio Estevez. Charlie Sheen is the sibling of Emilio Estevez. Janet Sheen is the mother of Emilio Estevez. Martin Sheen’s child is Ramón Estévez. Joe Estevez’s family name is Estevez. Martin Sheen’s family name is Estévez. Emilio Estevez is described by the source Obalky knih.cz. Ramón Estévez is a television actor. Emilio Estevez is a film director. Emilio Estevez is a citizen of the United States of America. Emilio Estevez is a writer. Estévez is said to be the same as Estevez. Emilio Estevez has a sibling named Renée Estévez. Emilio Estevez was born in New York City. Ramón Estévez is an actor. Charlie Sheen’s family name is Estévez. |
EFSum Emilio Estevez’s father is Martin Sheen. He was born in New York City in 1962 and is known for his work as an actor, film director, and writer. He has siblings named Ramón Estévez, Joe Estevez, Renée Estévez, and Charlie Sheen. Martin Sheen is the father of both Emilio Estevez and Ramón Estévez. |
In Table 4, we present an example of verbalized facts generated by each method. The example is from Mintaka. EFSum emphasizes evidence necessary to answer the question (i.e., Emilio Estevez’s father is Martin Sheen.) at the beginning, while excluding irrelevant details for brevity, as opposed to other baselines. Duplicated relations (i.e., relations of siblings) are aggregated, and it is clear that our method is more compact than KAPING and Rewrite. Comparing to KG2Text, it also sums up the given triples, but reduces a loss of essential information.